\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{100}

\bibitem{Kaluza:1921tu}
T.~{Kaluza}, \emph{{On the Unification Problem in Physics}},
  \href{https://doi.org/10.1142/S0218271818700017}{\emph{Int. J. Mod. Phys. D}
  {\bfseries 27} (2018) 1870001},
  [\href{https://arxiv.org/abs/1803.08616}{{\ttfamily 1803.08616}}].

\bibitem{Klein:1926fj}
O.~Klein, \emph{{The Atomicity of Electricity as a Quantum Theory Law}},
  \href{https://doi.org/10.1038/118516a0}{\emph{Nature} {\bfseries 118} (1926)
  516}.

\bibitem{Dienes:1996du}
K.~R. Dienes, \emph{{String theory and the path to unification: A Review of
  recent developments}},
  \href{https://doi.org/10.1016/S0370-1573(97)00009-4}{\emph{Phys. Rept.}
  {\bfseries 287} (1997) 447--525},
  [\href{https://arxiv.org/abs/hep-th/9602045}{{\ttfamily hep-th/9602045}}].

\bibitem{Green:2012oqa}
M.~B. Green, J.~H. Schwarz and E.~Witten, \emph{{Superstring Theory Vol. 1}}.
\newblock Cambridge Monographs on Mathematical Physics. Cambridge University
  Press, 2012,
  \href{https://doi.org/10.1017/CBO9781139248563}{10.1017/CBO9781139248563}.

\bibitem{Svrcek:2006yi}
P.~Svrcek and E.~Witten, \emph{{Axions In String Theory}},
  \href{https://doi.org/10.1088/1126-6708/2006/06/051}{\emph{JHEP} {\bfseries
  06} (2006) 051}, [\href{https://arxiv.org/abs/hep-th/0605206}{{\ttfamily
  hep-th/0605206}}].

\bibitem{Arvanitaki:2009fg}
A.~Arvanitaki, S.~Dimopoulos, S.~Dubovsky, N.~Kaloper and J.~March-Russell,
  \emph{{String Axiverse}},
  \href{https://doi.org/10.1103/PhysRevD.81.123530}{\emph{Phys. Rev.}
  {\bfseries D81} (2010) 123530},
  [\href{https://arxiv.org/abs/0905.4720}{{\ttfamily 0905.4720}}].

\bibitem{Hlozek:2014lca}
R.~Hlozek, D.~Grin, D.~J.~E. Marsh and P.~G. Ferreira, \emph{{A search for
  ultralight axions using precision cosmological data}},
  \href{https://doi.org/10.1103/PhysRevD.91.103512}{\emph{Phys. Rev.}
  {\bfseries D91} (2015) 103512},
  [\href{https://arxiv.org/abs/1410.2896}{{\ttfamily 1410.2896}}].

\bibitem{Marsh:2015xka}
D.~J.~E. Marsh, \emph{{Axion Cosmology}},
  \href{https://doi.org/10.1016/j.physrep.2016.06.005}{\emph{Phys. Rept.}
  {\bfseries 643} (2016) 1--79},
  [\href{https://arxiv.org/abs/1510.07633}{{\ttfamily 1510.07633}}].

\bibitem{Visinelli:2018utg}
L.~Visinelli and S.~Vagnozzi, \emph{{Cosmological window onto the string
  axiverse and the supersymmetry breaking scale}},
  \href{https://doi.org/10.1103/PhysRevD.99.063517}{\emph{Phys. Rev.}
  {\bfseries D99} (2019) 063517},
  [\href{https://arxiv.org/abs/1809.06382}{{\ttfamily 1809.06382}}].

\bibitem{Kinney:2018nny}
W.~H. Kinney, S.~Vagnozzi and L.~Visinelli, \emph{{The zoo plot meets the
  swampland: mutual (in)consistency of single-field inflation, string
  conjectures, and cosmological data}},
  \href{https://doi.org/10.1088/1361-6382/ab1d87}{\emph{Class. Quant. Grav.}
  {\bfseries 36} (2019) 117001},
  [\href{https://arxiv.org/abs/1808.06424}{{\ttfamily 1808.06424}}].

\bibitem{Randall:1999ee}
L.~Randall and R.~Sundrum, \emph{{A Large mass hierarchy from a small extra
  dimension}}, \href{https://doi.org/10.1103/PhysRevLett.83.3370}{\emph{Phys.
  Rev. Lett.} {\bfseries 83} (1999) 3370--3373},
  [\href{https://arxiv.org/abs/hep-ph/9905221}{{\ttfamily hep-ph/9905221}}].

\bibitem{Randall:1999vf}
L.~Randall and R.~Sundrum, \emph{{An Alternative to compactification}},
  \href{https://doi.org/10.1103/PhysRevLett.83.4690}{\emph{Phys. Rev. Lett.}
  {\bfseries 83} (1999) 4690--4693},
  [\href{https://arxiv.org/abs/hep-th/9906064}{{\ttfamily hep-th/9906064}}].

\bibitem{Binetruy:1999ut}
P.~Binetruy, C.~Deffayet and D.~Langlois, \emph{{Nonconventional cosmology from
  a brane universe}},
  \href{https://doi.org/10.1016/S0550-3213(99)00696-3}{\emph{Nucl. Phys.}
  {\bfseries B565} (2000) 269--287},
  [\href{https://arxiv.org/abs/hep-th/9905012}{{\ttfamily hep-th/9905012}}].

\bibitem{Binetruy:1999hy}
P.~Binetruy, C.~Deffayet, U.~Ellwanger and D.~Langlois, \emph{{Brane
  cosmological evolution in a bulk with cosmological constant}},
  \href{https://doi.org/10.1016/S0370-2693(00)00204-5}{\emph{Phys. Lett.}
  {\bfseries B477} (2000) 285--291},
  [\href{https://arxiv.org/abs/hep-th/9910219}{{\ttfamily hep-th/9910219}}].

\bibitem{Maartens:2010ar}
R.~Maartens and K.~Koyama, \emph{{Brane-World Gravity}},
  \href{https://doi.org/10.12942/lrr-2010-5}{\emph{Living Rev. Rel.} {\bfseries
  13} (2010) 5}, [\href{https://arxiv.org/abs/1004.3962}{{\ttfamily
  1004.3962}}].

\bibitem{Antoniadis:1998ig}
I.~Antoniadis, N.~Arkani-Hamed, S.~Dimopoulos and G.~R. Dvali, \emph{{New
  dimensions at a millimeter to a Fermi and superstrings at a TeV}},
  \href{https://doi.org/10.1016/S0370-2693(98)00860-0}{\emph{Phys. Lett.}
  {\bfseries B436} (1998) 257--263},
  [\href{https://arxiv.org/abs/hep-ph/9804398}{{\ttfamily hep-ph/9804398}}].

\bibitem{ArkaniHamed:1998nn}
N.~Arkani-Hamed, S.~Dimopoulos and G.~R. Dvali, \emph{{Phenomenology,
  astrophysics and cosmology of theories with submillimeter dimensions and TeV
  scale quantum gravity}},
  \href{https://doi.org/10.1103/PhysRevD.59.086004}{\emph{Phys. Rev.}
  {\bfseries D59} (1999) 086004},
  [\href{https://arxiv.org/abs/hep-ph/9807344}{{\ttfamily hep-ph/9807344}}].

\bibitem{Fitzpatrick:2007qr}
A.~L. Fitzpatrick, J.~Kaplan, L.~Randall and L.-T. Wang, \emph{{Searching for
  the Kaluza-Klein Graviton in Bulk RS Models}},
  \href{https://doi.org/10.1088/1126-6708/2007/09/013}{\emph{JHEP} {\bfseries
  09} (2007) 013}, [\href{https://arxiv.org/abs/hep-ph/0701150}{{\ttfamily
  hep-ph/0701150}}].

\bibitem{Johannsen:2008aa}
T.~Johannsen, \emph{{Constraints on the Size of Extra Dimensions from the
  Orbital Evolution of the Black-Hole X-Ray Binary XTE J1118+480}},
  \href{https://doi.org/10.1051/0004-6361/200912803}{\emph{Astron. Astrophys.}
  {\bfseries 507} (2009) 617},
  [\href{https://arxiv.org/abs/0812.0809}{{\ttfamily 0812.0809}}].

\bibitem{Salumbides:2015qwa}
E.~J. Salumbides, A.~N. Schellekens, B.~Gato-Rivera and W.~Ubachs,
  \emph{{Constraints on extra dimensions from precision molecular
  spectroscopy}},
  \href{https://doi.org/10.1088/1367-2630/17/3/033015}{\emph{New J. Phys.}
  {\bfseries 17} (2015) 033015},
  [\href{https://arxiv.org/abs/1502.02838}{{\ttfamily 1502.02838}}].

\bibitem{Gangopadhyay:2016qqa}
M.~R. Gangopadhyay and G.~J. Mathews, \emph{{Constraints on Brane-World
  Inflation from the CMB Power Spectrum: Revisited}},
  \href{https://doi.org/10.1088/1475-7516/2018/03/028}{\emph{JCAP} {\bfseries
  1803} (2018) 028}, [\href{https://arxiv.org/abs/1611.05123}{{\ttfamily
  1611.05123}}].

\bibitem{Pardo:2018ipy}
K.~Pardo, M.~Fishbach, D.~E. Holz and D.~N. Spergel, \emph{{Limits on the
  number of spacetime dimensions from GW170817}},
  \href{https://doi.org/10.1088/1475-7516/2018/07/048}{\emph{JCAP} {\bfseries
  1807} (2018) 048}, [\href{https://arxiv.org/abs/1801.08160}{{\ttfamily
  1801.08160}}].

\bibitem{Chakraborty:2016lxo}
S.~Chakraborty and S.~SenGupta, \emph{{Strong gravitational lensing --- A probe
  for extra dimensions and Kalb-Ramond field}},
  \href{https://doi.org/10.1088/1475-7516/2017/07/045}{\emph{JCAP} {\bfseries
  1707} (2017) 045}, [\href{https://arxiv.org/abs/1611.06936}{{\ttfamily
  1611.06936}}].

\bibitem{Chakravarti:2019aup}
K.~Chakravarti, S.~Chakraborty, K.~S. Phukon, S.~Bose and S.~SenGupta,
  \emph{{Constraining extra-spatial dimensions with observations of GW170817}},
   \href{https://arxiv.org/abs/1903.10159}{{\ttfamily 1903.10159}}.

\bibitem{Banerjee:2019sae}
I.~Banerjee, S.~Chakraborty and S.~SenGupta, \emph{{Decoding signatures of
  extra dimensions and estimating spin of quasars from the continuum
  spectrum}},  \href{https://arxiv.org/abs/1905.08043}{{\ttfamily 1905.08043}}.

\bibitem{Long:2002wn}
J.~C. Long, H.~W. Chan, A.~B. Churnside, E.~A. Gulbis, M.~C.~M. Varney and
  J.~C. Price, \emph{{Upper limits to submillimeter-range forces from extra
  space-time dimensions}},
  \href{https://arxiv.org/abs/hep-ph/0210004}{{\ttfamily hep-ph/0210004}}.

\bibitem{Visinelli:2017bny}
L.~Visinelli, N.~Bolis and S.~Vagnozzi, \emph{{Brane-world extra dimensions in
  light of GW170817}},
  \href{https://doi.org/10.1103/PhysRevD.97.064039}{\emph{Phys. Rev.}
  {\bfseries D97} (2018) 064039},
  [\href{https://arxiv.org/abs/1711.06628}{{\ttfamily 1711.06628}}].

\bibitem{Monitor:2017mdv}
{\scshape LIGO Scientific, Virgo, Fermi-GBM, INTEGRAL} collaboration, B.~P.
  Abbott et~al., \emph{{Gravitational Waves and Gamma-rays from a Binary
  Neutron Star Merger: GW170817 and GRB 170817A}},
  \href{https://doi.org/10.3847/2041-8213/aa920c}{\emph{Astrophys. J.}
  {\bfseries 848} (2017) L13},
  [\href{https://arxiv.org/abs/1710.05834}{{\ttfamily 1710.05834}}].

\bibitem{TheLIGOScientific:2017qsa}
{\scshape LIGO Scientific, Virgo} collaboration, B.~P. Abbott et~al.,
  \emph{{GW170817: Observation of Gravitational Waves from a Binary Neutron
  Star Inspiral}},
  \href{https://doi.org/10.1103/PhysRevLett.119.161101}{\emph{Phys. Rev. Lett.}
  {\bfseries 119} (2017) 161101},
  [\href{https://arxiv.org/abs/1710.05832}{{\ttfamily 1710.05832}}].

\bibitem{Callin:2004py}
P.~Callin and F.~Ravndal, \emph{{Higher order corrections to the Newtonian
  potential in the Randall-Sundrum model}},
  \href{https://doi.org/10.1103/PhysRevD.70.104009}{\emph{Phys. Rev.}
  {\bfseries D70} (2004) 104009},
  [\href{https://arxiv.org/abs/hep-ph/0403302}{{\ttfamily hep-ph/0403302}}].

\bibitem{Callin:2004bm}
P.~Callin, \emph{{Corrections to the Newtonian potential in the two-brane
  Randall-Sundrum model}},
  \href{https://arxiv.org/abs/hep-ph/0407054}{{\ttfamily hep-ph/0407054}}.

\bibitem{Callin:2004zm}
P.~Callin and F.~Ravndal, \emph{{Lagrangian formalism of gravity in the
  Randall-Sundrum model}},
  \href{https://doi.org/10.1103/PhysRevD.72.064026}{\emph{Phys. Rev.}
  {\bfseries D72} (2005) 064026},
  [\href{https://arxiv.org/abs/hep-ph/0412109}{{\ttfamily hep-ph/0412109}}].

\bibitem{Israel:1967wq}
W.~Israel, \emph{{Event horizons in static vacuum space-times}},
  \href{https://doi.org/10.1103/PhysRev.164.1776}{\emph{Phys. Rev.} {\bfseries
  164} (1967) 1776--1779}.

\bibitem{Israel:1967za}
W.~Israel, \emph{{Event horizons in static electrovac space-times}},
  \href{https://doi.org/10.1007/BF01645859}{\emph{Commun. Math. Phys.}
  {\bfseries 8} (1968) 245--260}.

\bibitem{Carter:1971zc}
B.~Carter, \emph{{Axisymmetric Black Hole Has Only Two Degrees of Freedom}},
  \href{https://doi.org/10.1103/PhysRevLett.26.331}{\emph{Phys. Rev. Lett.}
  {\bfseries 26} (1971) 331--333}.

\bibitem{Chrusciel:2012jk}
P.~T. Chrusciel, J.~Lopes~Costa and M.~Heusler, \emph{{Stationary Black Holes:
  Uniqueness and Beyond}},
  \href{https://doi.org/10.12942/lrr-2012-7}{\emph{Living Rev. Rel.} {\bfseries
  15} (2012) 7}, [\href{https://arxiv.org/abs/1205.6112}{{\ttfamily
  1205.6112}}].

\bibitem{Misner:1974qy}
C.~W. Misner, K.~S. Thorne and J.~A. Wheeler, \emph{{Gravitation}}.
\newblock W. H. Freeman, San Francisco, 1973.

\bibitem{Johannsen:2010xs}
T.~Johannsen and D.~Psaltis, \emph{{Testing the No-Hair Theorem with
  Observations in the Electromagnetic Spectrum: I. Properties of a Quasi-Kerr
  Spacetime}},
  \href{https://doi.org/10.1088/0004-637X/716/1/187}{\emph{Astrophys. J.}
  {\bfseries 716} (2010) 187--197},
  [\href{https://arxiv.org/abs/1003.3415}{{\ttfamily 1003.3415}}].

\bibitem{Johannsen:2010ru}
T.~Johannsen and D.~Psaltis, \emph{{Testing the No-Hair Theorem with
  Observations in the Electromagnetic Spectrum: II. Black-Hole Images}},
  \href{https://doi.org/10.1088/0004-637X/718/1/446}{\emph{Astrophys. J.}
  {\bfseries 718} (2010) 446--454},
  [\href{https://arxiv.org/abs/1005.1931}{{\ttfamily 1005.1931}}].

\bibitem{Luminet:1979nyg}
J.~P. Luminet, \emph{{Image of a spherical black hole with thin accretion
  disk}}, {\emph{Astron. Astrophys.} {\bfseries 75} (1979) 228--235}.

\bibitem{Falcke:1999pj}
H.~Falcke, F.~Melia and E.~Agol, \emph{{Viewing the shadow of the black hole at
  the galactic center}}, \href{https://doi.org/10.1086/312423}{\emph{Astrophys.
  J.} {\bfseries 528} (2000) L13},
  [\href{https://arxiv.org/abs/astro-ph/9912263}{{\ttfamily
  astro-ph/9912263}}].

\bibitem{Gralla:2019xty}
S.~E. Gralla, D.~E. Holz and R.~M. Wald, \emph{{Black Hole Shadows, Photon
  Rings, and Lensing Rings}},
  \href{https://arxiv.org/abs/1906.00873}{{\ttfamily 1906.00873}}.

\bibitem{Fish:2016jil}
{\scshape Event Horizon Telescope} collaboration, V.~L. Fish, K.~Akiyama, K.~L.
  Bouman, A.~A. Chael, M.~D. Johnson, S.~S. Doeleman et~al.,
  \emph{{Observing---and Imaging---Active Galactic Nuclei with the Event
  Horizon Telescope}},
  \href{https://doi.org/10.3390/galaxies4040054}{\emph{Galaxies} {\bfseries 4}
  (2016) 54}, [\href{https://arxiv.org/abs/1607.03034}{{\ttfamily
  1607.03034}}].

\bibitem{Akiyama:2019cqa}
{\scshape Event Horizon Telescope} collaboration, K.~Akiyama et~al.,
  \emph{{First M87 Event Horizon Telescope Results. I. The Shadow of the
  Supermassive Black Hole}},
  \href{https://doi.org/10.3847/2041-8213/ab0ec7}{\emph{Astrophys. J.}
  {\bfseries 875} (2019) L1}.

\bibitem{Akiyama:2019brx}
{\scshape Event Horizon Telescope} collaboration, K.~Akiyama et~al.,
  \emph{{First M87 Event Horizon Telescope Results. II. Array and
  Instrumentation}},
  \href{https://doi.org/10.3847/2041-8213/ab0c96}{\emph{Astrophys. J.}
  {\bfseries 875} (2019) L2}.

\bibitem{Akiyama:2019sww}
{\scshape Event Horizon Telescope} collaboration, K.~Akiyama et~al.,
  \emph{{First M87 Event Horizon Telescope Results. III. Data Processing and
  Calibration}},
  \href{https://doi.org/10.3847/2041-8213/ab0c57}{\emph{Astrophys. J.}
  {\bfseries 875} (2019) L3}.

\bibitem{Akiyama:2019bqs}
{\scshape Event Horizon Telescope} collaboration, K.~Akiyama et~al.,
  \emph{{First M87 Event Horizon Telescope Results. IV. Imaging the Central
  Supermassive Black Hole}},
  \href{https://doi.org/10.3847/2041-8213/ab0e85}{\emph{Astrophys. J.}
  {\bfseries 875} (2019) L4}.

\bibitem{Akiyama:2019fyp}
{\scshape Event Horizon Telescope} collaboration, K.~Akiyama et~al.,
  \emph{{First M87 Event Horizon Telescope Results. V. Physical Origin of the
  Asymmetric Ring}},
  \href{https://doi.org/10.3847/2041-8213/ab0f43}{\emph{Astrophys. J.}
  {\bfseries 875} (2019) L5}.

\bibitem{Akiyama:2019eap}
{\scshape Event Horizon Telescope} collaboration, K.~Akiyama et~al.,
  \emph{{First M87 Event Horizon Telescope Results. VI. The Shadow and Mass of
  the Central Black Hole}},
  \href{https://doi.org/10.3847/2041-8213/ab1141}{\emph{Astrophys. J.}
  {\bfseries 875} (2019) L6}.

\bibitem{Kerr:1963ud}
R.~P. Kerr, \emph{{Gravitational field of a spinning mass as an example of
  algebraically special metrics}},
  \href{https://doi.org/10.1103/PhysRevLett.11.237}{\emph{Phys. Rev. Lett.}
  {\bfseries 11} (1963) 237--238}.

\bibitem{Hernandez:1967zza}
W.~C. Hernandez, \emph{{Material Sources for the Kerr Metric}},
  \href{https://doi.org/10.1103/PhysRev.159.1070}{\emph{Phys. Rev.} {\bfseries
  159} (1967) 1070--1072}.

\bibitem{Argyres:1998qn}
P.~C. Argyres, S.~Dimopoulos and J.~March-Russell, \emph{{Black holes and
  submillimeter dimensions}},
  \href{https://doi.org/10.1016/S0370-2693(98)01184-8}{\emph{Phys. Lett.}
  {\bfseries B441} (1998) 96--104},
  [\href{https://arxiv.org/abs/hep-th/9808138}{{\ttfamily hep-th/9808138}}].

\bibitem{Chamblin:1999by}
A.~Chamblin, S.~W. Hawking and H.~S. Reall, \emph{{Brane world black holes}},
  \href{https://doi.org/10.1103/PhysRevD.61.065007}{\emph{Phys. Rev.}
  {\bfseries D61} (2000) 065007},
  [\href{https://arxiv.org/abs/hep-th/9909205}{{\ttfamily hep-th/9909205}}].

\bibitem{Kraus:1999it}
P.~Kraus, \emph{{Dynamics of anti-de Sitter domain walls}},
  \href{https://doi.org/10.1088/1126-6708/1999/12/011}{\emph{JHEP} {\bfseries
  12} (1999) 011}, [\href{https://arxiv.org/abs/hep-th/9910149}{{\ttfamily
  hep-th/9910149}}].

\bibitem{Emparan:1999wa}
R.~Emparan, G.~T. Horowitz and R.~C. Myers, \emph{{Exact description of black
  holes on branes}},
  \href{https://doi.org/10.1088/1126-6708/2000/01/007}{\emph{JHEP} {\bfseries
  01} (2000) 007}, [\href{https://arxiv.org/abs/hep-th/9911043}{{\ttfamily
  hep-th/9911043}}].

\bibitem{Garriga:1999yh}
J.~Garriga and T.~Tanaka, \emph{{Gravity in the brane world}},
  \href{https://doi.org/10.1103/PhysRevLett.84.2778}{\emph{Phys. Rev. Lett.}
  {\bfseries 84} (2000) 2778--2781},
  [\href{https://arxiv.org/abs/hep-th/9911055}{{\ttfamily hep-th/9911055}}].

\bibitem{Csaki:2000fc}
C.~Csaki, J.~Erlich, T.~J. Hollowood and Y.~Shirman, \emph{{Universal aspects
  of gravity localized on thick branes}},
  \href{https://doi.org/10.1016/S0550-3213(00)00271-6}{\emph{Nucl. Phys.}
  {\bfseries B581} (2000) 309--338},
  [\href{https://arxiv.org/abs/hep-th/0001033}{{\ttfamily hep-th/0001033}}].

\bibitem{Chamblin:2000md}
A.~Chamblin, C.~Csaki, J.~Erlich and T.~J. Hollowood, \emph{{Black diamonds at
  brane junctions}},
  \href{https://doi.org/10.1103/PhysRevD.62.044012}{\emph{Phys. Rev.}
  {\bfseries D62} (2000) 044012},
  [\href{https://arxiv.org/abs/hep-th/0002076}{{\ttfamily hep-th/0002076}}].

\bibitem{Dadhich:2000am}
N.~Dadhich, R.~Maartens, P.~Papadopoulos and V.~Rezania, \emph{{Black holes on
  the brane}}, \href{https://doi.org/10.1016/S0370-2693(00)00798-X}{\emph{Phys.
  Lett.} {\bfseries B487} (2000) 1--6},
  [\href{https://arxiv.org/abs/hep-th/0003061}{{\ttfamily hep-th/0003061}}].

\bibitem{Nojiri:2000yr}
S.~Nojiri, O.~Obregon, S.~D. Odintsov and S.~Ogushi, \emph{{Dilatonic brane
  world black holes, gravity localization and Newton constant}},
  \href{https://doi.org/10.1103/PhysRevD.62.064017}{\emph{Phys. Rev.}
  {\bfseries D62} (2000) 064017},
  [\href{https://arxiv.org/abs/hep-th/0003148}{{\ttfamily hep-th/0003148}}].

\bibitem{Casadio:2000py}
R.~Casadio and B.~Harms, \emph{{Black hole evaporation and large extra
  dimensions}},
  \href{https://doi.org/10.1016/S0370-2693(00)00840-6}{\emph{Phys. Lett.}
  {\bfseries B487} (2000) 209--214},
  [\href{https://arxiv.org/abs/hep-th/0004004}{{\ttfamily hep-th/0004004}}].

\bibitem{Giddings:2000mu}
S.~B. Giddings, E.~Katz and L.~Randall, \emph{{Linearized gravity in brane
  backgrounds}},
  \href{https://doi.org/10.1088/1126-6708/2000/03/023}{\emph{JHEP} {\bfseries
  03} (2000) 023}, [\href{https://arxiv.org/abs/hep-th/0002091}{{\ttfamily
  hep-th/0002091}}].

\bibitem{Aliev:2005bi}
A.~N. Aliev and A.~E. Gumrukcuoglu, \emph{{Charged rotating black holes on a
  3-brane}}, \href{https://doi.org/10.1103/PhysRevD.71.104027}{\emph{Phys.
  Rev.} {\bfseries D71} (2005) 104027},
  [\href{https://arxiv.org/abs/hep-th/0502223}{{\ttfamily hep-th/0502223}}].

\bibitem{Modgil:2001hm}
M.~S. Modgil, S.~Panda and G.~Sengupta, \emph{{Rotating brane world black
  holes}}, \href{https://doi.org/10.1142/S0217732302007442}{\emph{Mod. Phys.
  Lett.} {\bfseries A17} (2002) 1479--1488},
  [\href{https://arxiv.org/abs/hep-th/0104122}{{\ttfamily hep-th/0104122}}].

\bibitem{daRocha:2005xk}
R.~da~Rocha and C.~H. Coimbra-Araujo, \emph{{Variation in the luminosity of
  kerr quasars due to extra dimension in brane randall-sundrum model}},
  \href{https://doi.org/10.1088/1475-7516/2005/12/009}{\emph{JCAP} {\bfseries
  0512} (2005) 009}, [\href{https://arxiv.org/abs/astro-ph/0510318}{{\ttfamily
  astro-ph/0510318}}].

\bibitem{Aliev:2009cg}
A.~N. Aliev and P.~Talazan, \emph{{Gravitational Effects of Rotating Braneworld
  Black Holes}}, \href{https://doi.org/10.1103/PhysRevD.80.044023}{\emph{Phys.
  Rev.} {\bfseries D80} (2009) 044023},
  [\href{https://arxiv.org/abs/0906.1465}{{\ttfamily 0906.1465}}].

\bibitem{Amarilla:2011fx}
L.~Amarilla and E.~F. Eiroa, \emph{{Shadow of a rotating braneworld black
  hole}}, \href{https://doi.org/10.1103/PhysRevD.85.064019}{\emph{Phys. Rev.}
  {\bfseries D85} (2012) 064019},
  [\href{https://arxiv.org/abs/1112.6349}{{\ttfamily 1112.6349}}].

\bibitem{Aliev:2012rj}
A.~N. Aliev, G.~D. Esmer and P.~Talazan, \emph{{Strong Gravity Effects of
  Rotating Black Holes: Quasiperiodic Oscillations}},
  \href{https://doi.org/10.1088/0264-9381/30/4/045010}{\emph{Class. Quant.
  Grav.} {\bfseries 30} (2013) 045010},
  [\href{https://arxiv.org/abs/1205.2838}{{\ttfamily 1205.2838}}].

\bibitem{Neves:2012it}
J.~C.~S. Neves and C.~Molina, \emph{{Rotating black holes in a Randall-Sundrum
  brane with a cosmological constant}},
  \href{https://doi.org/10.1103/PhysRevD.86.124047}{\emph{Phys. Rev.}
  {\bfseries D86} (2012) 124047},
  [\href{https://arxiv.org/abs/1211.2848}{{\ttfamily 1211.2848}}].

\bibitem{Eiroa:2017uuq}
E.~F. Eiroa and C.~M. Sendra, \emph{{Shadow cast by rotating braneworld black
  holes with a cosmological constant}},
  \href{https://doi.org/10.1140/epjc/s10052-018-5586-6}{\emph{Eur. Phys. J.}
  {\bfseries C78} (2018) 91},
  [\href{https://arxiv.org/abs/1711.08380}{{\ttfamily 1711.08380}}].

\bibitem{Banerjee:2019cjk}
I.~Banerjee, B.~Mandal and S.~SenGupta, \emph{{Does black hole continuum
  spectrum signal higher curvature gravity in higher dimensions?}},
  \href{https://arxiv.org/abs/1905.12820}{{\ttfamily 1905.12820}}.

\bibitem{Schee:2008kz}
J.~Schee and Z.~Stuchlík, \emph{{Optical phenomena in the field of braneworld
  Kerr black holes}},
  \href{https://doi.org/10.1142/S0218271809014881}{\emph{Int. J. Mod. Phys.}
  {\bfseries D18} (2009) 983--1024},
  [\href{https://arxiv.org/abs/0810.4445}{{\ttfamily 0810.4445}}].

\bibitem{Stuchlik:2018qyz}
Z.~Stuchlík, D.~Charbulák and J.~Schee, \emph{{Light escape cones in local
  reference frames of Kerr–de Sitter black hole spacetimes and related black
  hole shadows}},
  \href{https://doi.org/10.1140/epjc/s10052-018-5578-6}{\emph{Eur. Phys. J.}
  {\bfseries C78} (2018) 180},
  [\href{https://arxiv.org/abs/1811.00072}{{\ttfamily 1811.00072}}].

\bibitem{Mertens:2016rhi}
F.~Mertens, A.~P. Lobanov, R.~C. Walker and P.~E. Hardee, \emph{{Kinematics of
  the jet in M 87 on scales of 100--1000 Schwarzschild radii}},
  \href{https://doi.org/10.1051/0004-6361/201628829}{\emph{Astron. Astrophys.}
  {\bfseries 595} (2016) A54},
  [\href{https://arxiv.org/abs/1608.05063}{{\ttfamily 1608.05063}}].

\bibitem{Blandford:1977ds}
R.~D. Blandford and R.~L. Znajek, \emph{{Electromagnetic extractions of energy
  from Kerr black holes}},
  \href{https://doi.org/10.1093/mnras/179.3.433}{\emph{Mon. Not. Roy. Astron.
  Soc.} {\bfseries 179} (1977) 433--456}.

\bibitem{Rieger:2012en}
F.~M. Rieger and F.~Aharonian, \emph{{Probing the central black hole in M87
  with gamma-rays}},
  \href{https://doi.org/10.1142/S0217732312300303}{\emph{Mod. Phys. Lett.}
  {\bfseries A27} (2012) 1230030},
  [\href{https://arxiv.org/abs/1208.2702}{{\ttfamily 1208.2702}}].

\bibitem{Giddings:2019jnp}
S.~B. Giddings, \emph{{Searching for quantum black hole structure with the
  Event Horizon Telescope}},
  \href{https://arxiv.org/abs/1904.05287}{{\ttfamily 1904.05287}}.

\bibitem{Moffat:2019uxp}
J.~W. Moffat and V.~T. Toth, \emph{{The masses and shadows of the black holes
  Sagittarius A* and M87 in modified gravity (MOG)}},
  \href{https://arxiv.org/abs/1904.04142}{{\ttfamily 1904.04142}}.

\bibitem{Nokhrina:2019sxv}
E.~E. Nokhrina, L.~I. Gurvits, V.~S. Beskin, M.~Nakamura, K.~Asada and K.~Hada,
  \emph{{M87 black hole mass and spin estimate through the position of the jet
  boundary shape break}},  \href{https://arxiv.org/abs/1904.05665}{{\ttfamily
  1904.05665}}.

\bibitem{Abdikamalov:2019ztb}
A.~B. Abdikamalov, A.~A. Abdujabbarov, D.~Ayzenberg, D.~Malafarina, C.~Bambi
  and B.~Ahmedov, \emph{{A black hole mimicker hiding in the shadow: Optical
  properties of the $\gamma$ metric}},
  \href{https://arxiv.org/abs/1904.06207}{{\ttfamily 1904.06207}}.

\bibitem{Held:2019xde}
A.~Held, R.~Gold and A.~Eichhorn, \emph{{Asymptotic safety casts its shadow}},
  \href{https://arxiv.org/abs/1904.07133}{{\ttfamily 1904.07133}}.

\bibitem{Wei:2019pjf}
S.-W. Wei, Y.-C. Zou, Y.-X. Liu and R.~B. Mann, \emph{{Curvature radius and
  Kerr black hole shadow}},  \href{https://arxiv.org/abs/1904.07710}{{\ttfamily
  1904.07710}}.

\bibitem{Tamburini:2019vrf}
F.~Tamburini, B.~Thid{\'e} and M.~Della~Valle, \emph{{Measurement of the spin
  of the M87 black hole from its observed twisted light}},
  \href{https://arxiv.org/abs/1904.07923}{{\ttfamily 1904.07923}}.

\bibitem{Davoudiasl:2019nlo}
H.~Davoudiasl and P.~B. Denton, \emph{{Ultra Light Boson Dark Matter and Event
  Horizon Telescope Observations of M87*}},
  \href{https://arxiv.org/abs/1904.09242}{{\ttfamily 1904.09242}}.

\bibitem{Ovgun:2019yor}
A.~{\"O}vg{\"u}n, {\.I}.~Sakallı and H.~Mutuk, \emph{{Quasinormal modes of
  Schwarzschild Black Hole and Damour-Solodukhin Wormhole via Feedforward
  Neural Network Method}},  \href{https://arxiv.org/abs/1904.09509}{{\ttfamily
  1904.09509}}.

\bibitem{Cunha:2019dwb}
P.~V.~P. Cunha, C.~A.~R. Herdeiro and E.~Radu, \emph{{Spontaneously scalarised
  Kerr black holes}},  \href{https://arxiv.org/abs/1904.09997}{{\ttfamily
  1904.09997}}.

\bibitem{Wang:2019tjc}
M.~Wang, S.~Chen, J.~Wang and J.~Jing, \emph{{Shadow of a Schwarzschild black
  hole surrounded by a Bach-Weyl ring}},
  \href{https://arxiv.org/abs/1904.12423}{{\ttfamily 1904.12423}}.

\bibitem{Hui:2019aqm}
L.~Hui, D.~Kabat, X.~Li, L.~Santoni and S.~S.~C. Wong, \emph{{Black Hole Hair
  from Scalar Dark Matter}},
  \href{https://arxiv.org/abs/1904.12803}{{\ttfamily 1904.12803}}.

\bibitem{Bambi:2019tjh}
C.~Bambi, K.~Freese, S.~Vagnozzi and L.~Visinelli, \emph{{Testing the
  rotational nature of the supermassive object M87* from the circularity and
  size of its first image}},
  \href{https://arxiv.org/abs/1904.12983}{{\ttfamily 1904.12983}}.

\bibitem{Konoplya:2019sns}
R.~A. Konoplya, \emph{{Shadow of a black hole surrounded by dark matter}},
  \href{https://doi.org/10.1016/j.physletb.2019.05.043}{\emph{Phys. Lett.}
  {\bfseries B795} (2019) 1--6},
  [\href{https://arxiv.org/abs/1905.00064}{{\ttfamily 1905.00064}}].

\bibitem{Nemmen:2019idv}
R.~Nemmen, \emph{{The Spin of M87*}},
  \href{https://arxiv.org/abs/1905.02143}{{\ttfamily 1905.02143}}.

\bibitem{Chen:2019fsq}
Y.~Chen, J.~Shu, X.~Xue, Q.~Yuan and Y.~Zhao, \emph{{Probing Axions with Event
  Horizon Telescope Polarimetric Measurements}},
  \href{https://arxiv.org/abs/1905.02213}{{\ttfamily 1905.02213}}.

\bibitem{Gyulchev:2019tvk}
G.~Gyulchev, P.~Nedkova, T.~Vetsov and S.~Yazadjiev, \emph{{Image of the
  Janis-Newman-Winicour naked singularity with a thin accretion disk}},
  \href{https://arxiv.org/abs/1905.05273}{{\ttfamily 1905.05273}}.

\bibitem{Shaikh:2019jfr}
R.~Shaikh, P.~Banerjee, S.~Paul and T.~Sarkar, \emph{{Strong gravitational
  lensing by wormholes}},  \href{https://arxiv.org/abs/1905.06932}{{\ttfamily
  1905.06932}}.

\bibitem{Firouzjaee:2019aij}
J.~T. Firouzjaee and A.~Allahyari, \emph{{Black hole shadow with a cosmological
  constant for cosmological observers}},
  \href{https://arxiv.org/abs/1905.07378}{{\ttfamily 1905.07378}}.

\bibitem{Konoplya:2019nzp}
R.~A. Konoplya, C.~Posada, Z.~Stuchl{\'\i}k and A.~Zhidenko, \emph{{Stable
  Schwarzschild stars as black-hole mimickers}},
  \href{https://arxiv.org/abs/1905.08097}{{\ttfamily 1905.08097}}.

\bibitem{Kokubu:2019jdx}
T.~Kokubu and T.~Harada, \emph{{Burst particle creation in gravitational
  collapse to a horizonless compact object}},
  \href{https://arxiv.org/abs/1905.07981}{{\ttfamily 1905.07981}}.

\bibitem{Giddings:2019vvj}
S.~B. Giddings, \emph{{Black holes in the quantum universe}},
  \href{https://arxiv.org/abs/1905.08807}{{\ttfamily 1905.08807}}.

\bibitem{Kawashima:2019ljv}
T.~Kawashima, M.~Kino and K.~Akiyama, \emph{{Black Hole Spin Signature in the
  Black Hole Shadow of M87 in the Flaring State}},
  \href{https://arxiv.org/abs/1905.10717}{{\ttfamily 1905.10717}}.

\bibitem{Kumaran:2019qqp}
Y.~Kumaran and A.~{\"O}vg{\"u}n, \emph{{Weak Deflection Angle of Extended
  Uncertainty Principle Black Holes}},
  \href{https://arxiv.org/abs/1905.11710}{{\ttfamily 1905.11710}}.

\bibitem{Bar:2019pnz}
N.~Bar, K.~Blum, T.~Lacroix and P.~Panci, \emph{{Looking for ultralight dark
  matter near supermassive black holes}},
  \href{https://arxiv.org/abs/1905.11745}{{\ttfamily 1905.11745}}.

\bibitem{Jusufi:2019nrn}
K.~Jusufi, M.~Jamil, P.~Salucci, T.~Zhu and S.~Haroon, \emph{{Black Hole
  Surrounded by Dark Matter Halo in the M87 Galactic Center and its
  Identification with Shadow Images}},
  \href{https://arxiv.org/abs/1905.11803}{{\ttfamily 1905.11803}}.

\bibitem{Exirifard:2019ile}
Q.~Exirifard, \emph{{Gravitomagnetism in Modified theory of Gravity}},
  \href{https://arxiv.org/abs/1906.02013}{{\ttfamily 1906.02013}}.

\bibitem{Roy:2019esk}
R.~Roy and U.~A. Yajnik, \emph{{Evolution of black hole shadow in the presence
  of ultralight bosons}},  \href{https://arxiv.org/abs/1906.03190}{{\ttfamily
  1906.03190}}.

\bibitem{Bambi:2019xzp}
C.~Bambi, \emph{{Astrophysical Black Holes: A Review}},  2019,
  \href{https://arxiv.org/abs/1906.03871}{{\ttfamily 1906.03871}}.

\bibitem{Long:2019nox}
F.~Long, S.~Chen and J.~Jing, \emph{{Shadow of a rotating squashed Kaluza-Klein
  black hole}},  \href{https://arxiv.org/abs/1906.04456}{{\ttfamily
  1906.04456}}.

\end{thebibliography}\endgroup

\documentclass[aps, prd, twocolumn, preprintnumbers, superscriptaddress, nofootinbib, floatfix]{revtex4-1}

\usepackage{amsmath}
\usepackage{bm}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{epstopdf}
\usepackage{hyperref}
\usepackage{array}
\usepackage{soul}
\usepackage{color}
\usepackage[utf8]{inputenc}

\enlargethispage{\baselineskip}

\hypersetup{
     colorlinks   = true,
     citecolor    = blue
}

\begin{document}

\title{Hunting for extra dimensions in the shadow of M87*}

\newcommand{\OKCAFF}{\affiliation{The Oskar Klein Centre for Cosmoparticle Physics, Department of Physics, Stockholm University, \\AlbaNova Universitetscentrum, Roslagstullsbacken 21A, SE-106 91 Stockholm, Sweden}}
\newcommand{\FIRSTAFF}{\affiliation{Department of Physics and Astronomy, Uppsala University, L\"agerhyddsv\"agen 1, 75120 Uppsala, Sweden}}
\newcommand{\SECONDAFF}{\affiliation{Nordita, KTH Royal Institute of Technology and Stockholm University, Roslagstullsbacken 23, 10691 Stockholm, Sweden}}
\newcommand{\THIRDAFF}{\affiliation{Kavli Institute for Cosmology (KICC) and Institute of Astronomy,\\University of Cambridge, Madingley Road, Cambridge CB3 0HA, United Kingdom}}
\newcommand{\FOURTHAFF}{\affiliation{Gravitation Astroparticle Physics Amsterdam (GRAPPA), Institute for Theoretical Physics Amsterdam and\\Delta Institute for Theoretical Physics, University of Amsterdam, Science Park 904, 1098 XH Amsterdam, The Netherlands}}
\author{Sunny Vagnozzi}\email{sunny.vagnozzi@fysik.su.se} \OKCAFF \SECONDAFF \THIRDAFF
\author{Luca Visinelli}\email{luca.visinelli@physics.uu.se} \SECONDAFF \FIRSTAFF \FOURTHAFF
\date{\today}

\begin{abstract}
The Event Horizon Telescope has recently provided the first image of the dark shadow around the supermassive black hole M87*. The observation of a highly circular shadow provides strong limits on deviations of M87*'s quadrupole moment from the Kerr value. We show that the absence of such a deviation can be used to constrain the physics of extra dimensions of spacetime. Focusing on the Randall-Sundrum AdS$_5$ brane-world scenario, we show that the observation of M87*'s dark shadow sets the limit $\ell \lesssim 170\,{\rm AU}$, where $\ell$ is the AdS$_5$ curvature radius. This limit is among the first quantitative constraints on exotic physics obtained from the extraordinary first ever image of the dark shadow of a black hole.
\end{abstract}

\maketitle

\section{Introduction}
\label{sec:intro}

The possible existence of extra dimensions, first postulated by Kaluza~\cite{Kaluza:1921tu} and Klein~\cite{Klein:1926fj} in an attempt to unify gravity and electromagnetism, is among the most remarkable predictions of string theory and of $p$-branes~\cite{Dienes:1996du,Green:2012oqa}, albeit virtually inaccessible to current experimental tests.~\footnote{A related important testable prediction of string theory is the existence of a multitude of light pseudoscalar fields known as the axiverse (see e.g.~\cite{Svrcek:2006yi,Arvanitaki:2009fg,Hlozek:2014lca, Marsh:2015xka,Visinelli:2018utg} for works examining the observational testability of this prediction, and also~\cite{Kinney:2018nny} for further related work).} An alternative possibility which can be experimentally probed consists of a non-compactified, large extra dimension: such a scenario occurs for instance in the Randall-Sundrum (RS) model~\cite{Randall:1999ee,Randall:1999vf}. The RS scenario features a five-dimensional AdS space (AdS$_5$) brane-world model where the extra dimension has an infinite size and a negative bulk cosmological constant. The matter fields and the gauge fields of the electromagnetic, weak and strong forces are restricted to a three-dimensional brane, while gravity is free to propagate on the bulk of the AdS$_5$ spacetime. The tension on the brane is chosen so that General Relativity is recovered in the low-energy regime~\cite{Binetruy:1999ut,Binetruy:1999hy}. A review of brane-world gravity models is given in~\cite{Maartens:2010ar}.

In general, the existence of brane-world extra dimensions can be observationally tested through a variety of methods, which include astrophysical observations, precision tests of gravity on small scales, and collider searches~\cite{Antoniadis:1998ig,ArkaniHamed:1998nn,Fitzpatrick:2007qr,Johannsen:2008aa,
Salumbides:2015qwa,Gangopadhyay:2016qqa,Pardo:2018ipy,Chakraborty:2016lxo,Chakravarti:2019aup,Banerjee:2019sae}. In particular, these observations allow to constrain the AdS$_5$ radius of curvature $\ell$, which is a new parameter entering the metric of the space-time. Currently, the best limits on $\ell$ come from the lack of detection of macroscopic forces at laboratory distances, which set the constraint $\ell \lesssim 1\,$mm~\cite{Long:2002wn}. Since gravity and electromagnetism behave in a fundamentally different way in the RS model, studying the propagation of gravitational and electromagnetic signals from a common source is a powerful way to test the existence of brane-world extra dimensions. In a previous paper~\cite{Visinelli:2017bny}, we used the multi-messenger detection of the gravitational wave (GW) event GW170817 and its electromagnetic counterpart GRB170817A~\cite{Monitor:2017mdv,TheLIGOScientific:2017qsa} from a binary neutron star merger to test the RS scenario. The sole measurement of the time lag between the two multi-messenger counterparts allowed us to place the bound $\ell \lesssim {\cal O}({\rm Mpc})$.

Besides altering the propagation of GWs, brane-world extra dimensions are expected to affect the Newtonian potential generated by a massive object lying on the brane (see e.g.~\cite{Randall:1999ee,Randall:1999vf,Callin:2004py,Callin:2004bm,Callin:2004zm}). These modifications can be quite relevant for the metric describing the vicinity of black holes (BHs), since they would alter the BH quadrupole moment with respect to the expected value in the absence of extra dimensions. These modifications are particularly important since the no-hair theorem/conjecture states that BH solutions to the Einstein-Maxwell equations of GR and electromagnetism should be completely characterised by three parameters: mass, electric charge, and angular momentum~\cite{Israel:1967wq,Israel:1967za,Carter:1971zc, Chrusciel:2012jk,Misner:1974qy}. In particular, knowledge of these three quantities would be enough to uniquely determine the BH quadrupole moment: any deviation from such canonical value would be an indication that new physics is at play~\cite{Johannsen:2010xs,Johannsen:2010ru}.

In~\cite{Luminet:1979nyg}, it was shown that the combination of a BH event horizon and gravitational lensing of nearby photons leads to the appearance of a dark shadow, which should be detectable using very long baseline interferometry (VLBI)~\cite{Falcke:1999pj} (see also~\cite{Gralla:2019xty}). Recently, the Event Horizon Telescope (EHT) collaboration~\cite{Fish:2016jil} succeeded for the first time in providing a breathtaking image of the dark shadow of the supermassive black hole (SMBH) M87*, which resides at the center of the giant elliptical galaxy M87~\cite{Akiyama:2019cqa} (see also~\cite{Akiyama:2019brx,Akiyama:2019sww,Akiyama:2019bqs,Akiyama:2019fyp,Akiyama:2019eap}).

In the simplest scenario wherein M87* is a Kerr BH, and gravity behaves as per standard expectations, the dark shadow of M87* should be highly circular. This expectation is indeed met in the results of the EHT collaboration~\cite{Akiyama:2019cqa}, which reports deviations from circularity in the image of the dark shadow of M87* to be of order $10\%$ at most. Still in~\cite{Akiyama:2019cqa}, such a constraint was used to set an upper limit on the relative deviation of M87*'s quadrupole moment fromthe standard value $Q_{\rm Kerr}$: $|\Delta Q/Q_{\rm Kerr}| \equiv \epsilon \lesssim 4$. 

In this short paper, it is our goal to revisit the physics of brane-world extra dimensions in light of the highly circular dark shadow of M87* detected by the EHT collaboration. The rest of this paper is then organized as follows. In Sec.~\ref{sec:rotating}, we first revisit the computation of the quadrupole moment of Kerr BHs within the RS scenario, wherein a fifth dimension with curvature radius $\ell$ is present. Still in Sec.~\ref{sec:rotating}, we then estimate the relative deviation of the quadrupole moment compared to standard expectations as a function of $\ell$, $\epsilon(\ell)$. In Sec.~\ref{sec:results} we use the limit $\epsilon(\ell) \lesssim 4$ set by the EHT collaboration to set a new independent upper limit on the AdS$_5$ curvature radius $\ell$ within the Randall-Sundrum scenario. We provide closing remarks in Sec.~\ref{sec:conclusions}.

\section{Rotating black holes in 4 and 5 dimensions}
\label{sec:rotating}

Let us first consider a Kerr BH in 4 dimensions. The BH has mass $M$ and angular momentum ${\bf J}$, and we further define the rotational parameter $a \equiv |{\bf J}/M|$. Within the usual spherical coordinate system $(r, \theta, \varphi)$, with $r=\sqrt{x^2+y^2+z^2}$, where $(x,y,z)$ are the usual Cartesian coordinates, and in units where $c=G=1$, the square infinitesimal line element of a Kerr BH in 4 dimensions is given by~\cite{Kerr:1963ud}:
\begin{eqnarray}
ds^2 = &-& \left ( 1-\frac{2Mr}{\Sigma} \right )dt^2 + \frac{\Sigma}{\Delta}dr^2+\Sigma d\theta^2 \nonumber \\
&+& \left ( r^2+a^2+\frac{2Mra^2}{\Sigma}\sin^2\theta \right )\sin^2\theta d\varphi^2 \nonumber \\
&-& \frac{4Mra\sin^2\theta}{\Sigma}dtd\varphi\,.
\label{eq:kerr}
\end{eqnarray}
Let us further define the function $\gamma(r,z)$, defined from the real and positive solution of the equation $\gamma^4-(r^2-a^2)\gamma^2-a^2z^2 = 0$. Then, in the weak field limit, the Newtonian gravitational potential $\Phi(r,\theta)$ generated by a source of mass $M$ is given by~\cite{Hernandez:1967zza}:
\begin{eqnarray}
\Phi(r,\theta) = -\frac{M\gamma^3}{\gamma^4+a^2z^2}\,.
\label{eq:potential}
\end{eqnarray}
As shown in~\cite{Hernandez:1967zza}, the Newtonian gravitational potential can then be expanded in multipoles to yield (up to order $(a/r)^2$):
\begin{eqnarray}
\Phi(r,\theta) = -\frac{M}{r} \left [ 1+ \left ( \frac{a}{r} \right )^2{\cal P}_2(\cos\theta) +... \right ]\,,
\label{eq:multipoles}
\end{eqnarray}
where ${\cal P}_2(x) = (3x^2-1)/2$ is the second Legendre polynomial. As expected, for $a=0$ the potential is spherically symmetric, whereas the correction to spherical symmetry is given by the second term on the right-hand side of Eq.~(\ref{eq:multipoles}). In particular, from the correction we read off the Kerr BH quadrupole, given by $Q_{\rm Kerr}=Ma^2$.

We now need to correct Eq.~(\ref{eq:multipoles}) in the presence of a fifth dimension. The process of matter collapse to a BH within the RS model was studied in a series of seminal works~\cite{Argyres:1998qn,Chamblin:1999by,Kraus:1999it,Emparan:1999wa,Garriga:1999yh,Csaki:2000fc,Chamblin:2000md,
Dadhich:2000am,Nojiri:2000yr,Casadio:2000py}. In particular~\cite{Chamblin:1999by} studied the case of a black string living in the AdS$_5$ bulk, whose intersection with the 3-brane would appear as a BH to an observer living on the brane. However, such a ``projected" BH still carries information about the fifth dimension, as discussed in~\cite{Dadhich:2000am}. When projected onto the brane, gravitational field effects in the bulk effectively endow the BH with a \textit{tidal charge} $\beta$, carrying information about the extra dimension and in particular about the AdS$_5$ curvature radius $\ell$. In other words, the BH living on the brane carries an effective charge (which might be positive or negative) due to the tidal influence of the bulk.

These earlier works focused on brane-world BHs for non-rotating spacetimes. As shown in~\cite{Randall:1999ee,Randall:1999vf,Giddings:2000mu}, the continuous spectrum of Kaluza-Klein modes in RS models leads to a correction to the gravitational potential generated by a point source which at low energies goes as $\Delta \Phi \approx 2M\ell^2/(3r^3)$. What about rotating spacetimes, \textit{i.e.} brane-world Kerr BHs? The square infinitesimal line element of a Kerr BH in AdS$_5$ is given by~\cite{Aliev:2005bi}:~\footnote{See also~\cite{Modgil:2001hm,daRocha:2005xk,Aliev:2009cg,Amarilla:2011fx,Aliev:2012rj,
Neves:2012it,Eiroa:2017uuq,Banerjee:2019cjk,Schee:2008kz,Stuchlik:2018qyz} for important works on rotating BHs in brane-world models.}
\begin{eqnarray}
ds^2 = &-&\frac{\Delta}{\Sigma} \left (dt-a\sin^2\theta d\phi \right ) ^2 + \Sigma \left ( \frac{dr^2}{\Delta} + d\phi^2 \right ) \nonumber \\
&+& \frac{\sin^2\theta}{\Sigma} \left [ adt - \left ( r^2+a^2 \right )d\phi \right ]^2\,,
\label{eq:kerr5d}
\end{eqnarray}
where we have defined:
\begin{eqnarray}
\Delta \equiv r^2+a^2-2Mr+\beta\,, \quad \Sigma \equiv r^2+a^2\cos^2\theta\,,
\label{eq:deltasigma}
\end{eqnarray}
with $\beta$ the tidal charge of the BH, which records information from the bulk projected onto the brane. The metric in Eq.~(\ref{eq:kerr5d}) may look familiar: in fact, it is exactly equivalent to the Kerr-Newman metric describing a rotating charged BH in 4 dimensions, where this time the charge $Q$ has been replaced by the tidal charge $\beta$ in the definition of $\Delta$ in Eq.~(\ref{eq:deltasigma}). In fact, it is easy to see that for $\beta=0$ one recovers the Kerr metric of Eq.~(\ref{eq:kerr}). This result is totally analogous to the one found in~\cite{Dadhich:2000am}, where it was found that a Schwarzschild (\textit{i.e.} non-rotating, uncharged) brane-world BH is described by a Reissner-Nordstr\"{o}m (\textit{i.e.} non-rotating, charged) metric, with the BH charge replaced by the tidal charge.

One can show that the earlier estimate of $\Delta \Phi \approx 2M\ell^2/(3r^3)$ is valid to order $(a/r)^2$ (\textit{i.e.} the order relevant for computing the quadrupole) even when considering rotating spacetimes. In other words, the quadrupole does not receive corrections depending on the spin of the BH. This can be seen by following the approach of~\cite{daRocha:2005xk}, wherein one obtains that the leading order correction scales as $\ell^2a^2/r^4$, and thus does not affect the computation of the quadrupole. In summary, the Newtonian gravitational potential for a Kerr BH in the RS model is given, up to order $(a/r)^2$, by:
\begin{eqnarray}
\Phi(r,\theta) = -\frac{GM}{r} \left [ 1+ \left ( \frac{a}{r} \right )^2{\cal P}_2(\cos\theta)+\frac{2\ell^2}{3r^2} +... \right ].
\label{eq:multipolesbrane}
\end{eqnarray}

In all previous equations, $\theta$ is the BH observation angle, which corresponds to the angle between the line-of-sight and the spin axis. For the specific case of M87*, the jet inclination with respect to the line-of-sight is $\theta_{\rm jet} \simeq 17^{\circ}$~\cite{Mertens:2016rhi}, which coincides with the angle between the line-of-sight and the spin axis $\theta$ under the well-motivated assumption that the jet is powered by the spin of the BH and is approximately aligned with the spin axis (through the Blandford-Znajek mechanism~\cite{Blandford:1977ds} or variants thereof). Hereafter, we shall therefore set $\theta = \theta_{\rm jet} \simeq 17^{\circ}$. Notice that for such low observation angle, ${\cal P}_2(\cos\theta) \approx 1$, an approximation which for simplicity we adopt moving forward.~\footnote{There are in fact indications that $\theta$ might even be lower than $17^{\circ}$~\cite{Rieger:2012en}. This would make our approximation of ${\cal P}_2(\cos\theta) \approx 1$ even more accurate.}

\section{Results}
\label{sec:results}

Using the modified Newtonian potential in the presence of extra dimensions estimated in Eq.~(\ref{eq:multipolesbrane}), and using our approximation for ${\cal P}_2(\cos\theta) \approx 1$, we can read off the quadrupole moment $Q_{\rm Kerr,RS} \approx Ma^2+2M\ell^2/3$ and correspondingly the deviation from the Kerr value $\Delta Q \equiv Q_{\rm Kerr,RS}-Q_{\rm Kerr}$, with $Q_{\rm Kerr}=Ma^2$. We find that the deviation $\Delta Q$, and correspondingly the relative deviation $\epsilon$, are given by:
\begin{eqnarray}
\Delta Q &\approx& \frac{2M\ell^2}{3}\,, \label{eq:deltaq}\\
\epsilon &\equiv& \frac{\Delta Q}{Q_{\rm Kerr}} \approx \frac{2\ell^2}{3a^2}\,. \label{eq:epsilon}
\end{eqnarray}

Using our estimate for $\Delta Q$ together with the upper limit on $\epsilon$ reported by the EHT collaboration implies:
\begin{eqnarray}
\epsilon \approx \frac{2}{3}\frac{\ell^2}{a^2} \lesssim 4 \implies \frac{2}{3}\ell^2 \lesssim 4a^2\,.
\label{eq:upperlimit}
\end{eqnarray}
However, for a Kerr BH the angular momentum is bounded by the Kerr limit $a \leq M$, which implies that $4a^2 \lesssim R_s^2$, where $R_s=2M$ is the Schwarzschild radius for the object in question. Armed with this further inequality, we see that Eq.~(\ref{eq:upperlimit}) implies:
\begin{eqnarray}
\ell \lesssim \sqrt{\frac{3}{2}}R_s\,.
\label{eq:upperlimitnew}
\end{eqnarray}
In~\cite{Akiyama:2019eap}, the mass of M87* was estimated to be $M_{M87*} = (6.6\pm 0.7)\times10^{9}\,M_{\odot}$, in units of the mass of the Sun $M_{\odot} \approx 2\times 10^{30}\,$kg. Using this result, we find $R_s \approx 2.5\times 10^{13}\,{\rm m}$, which finally results in the following upper limit on AdS$_5$ curvature radius:
\begin{eqnarray}
\ell \lesssim 2.5\times 10^{13}\,{\rm m} \approx 170\,{\rm AU}\,,
\label{eq:upperlimitfinal}
\end{eqnarray}
with $1\,{\rm AU}$ being one astronomical unit.

The upper limit obtained in Eq.~(\ref{eq:upperlimitfinal}) is still not competitive with the ${\cal O}({\rm mm})$ constraint obtained in~\cite{Long:2002wn}, but represents a substantial improvement over the limit we obtained from GW170817 in~\cite{Visinelli:2017bny}. More importantly, it is an independent limit and among the first limits on exotic physics obtained by imaging the dark shadow of M87*.~\footnote{See also~\cite{Giddings:2019jnp,Moffat:2019uxp,Nokhrina:2019sxv,Abdikamalov:2019ztb,Held:2019xde,Wei:2019pjf,Tamburini:2019vrf,Davoudiasl:2019nlo, Ovgun:2019yor,Cunha:2019dwb,Wang:2019tjc,Hui:2019aqm,Bambi:2019tjh,Konoplya:2019sns,Nemmen:2019idv,Chen:2019fsq,Gyulchev:2019tvk,Shaikh:2019jfr, Firouzjaee:2019aij,Konoplya:2019nzp,Kokubu:2019jdx,Giddings:2019vvj,Kawashima:2019ljv,Kumaran:2019qqp,Bar:2019pnz,Jusufi:2019nrn,Exirifard:2019ile,
Roy:2019esk,Bambi:2019xzp,Long:2019nox} for other works exploring fundamental physics and properties of BHs or BH mimickers in light of the imaged dark shadow of M87* or future observations from the EHT.}

\section{Conclusions}
\label{sec:conclusions}

In this short paper we have shown how the image of the dark shadow of M87*, recently provided by the Event Horizon Telescope collaboration, allows us to probe the physics of extra dimensions. Focusing on the Randall-Sundrum brane-world scenario, wherein our Universe consists of a 3-brane within an AdS$_5$ space, we have set an upper limit on the AdS$_5$ curvature radius $\ell$, finding $\ell \lesssim 170\,{\rm AU}$. This upper limit is far from being competitive with current constraints on extra dimensions from precision tests of gravity on ${\cal O}({\rm mm})$ scale. While weak, we believe this constraint is nonetheless valuable because it is obtained in a completely independent way with respect to previous limits. Moreover, it serves as a proof-of-principle for the possibility of constraining fundamental physics already from the first image of the dark shadow of a black hole.

The coming years will certainly be an extremely exciting time as far as tests of General Relativity and alternative theories of gravity are concerned. The success of the Event Horizon Telescope in imaging the dark shadow of M87* pairs up with LIGO's success in detecting an ever-increasing number of GW events: together, the two provide us further tests of the behaviour of gravity on astronomical and cosmological scales. It is exciting to see how we can use available data and build upon the considerable efforts put into action to deploy the first ever image of a black hole shadow, to probe physics describing the very structure of our space-time, testing the possible existence of extra dimensions beyond those which our senses are able to experience. We look forward to improvements in VLBI technology, as well as the delivery of the first image of Sgr A*, which will allow us to refine tests of fundamental physics by imaging the dark shadows of astrophysical black holes.

\begin{acknowledgments}
We thank Cosimo Bambi, Katherine Freese, and Maurizio Giannotti for useful discussions on the subject. We acknowledge support by the Vetenskapsr\r{a}det (Swedish Research Council) through contract No. 638-2013-8993 and the Oskar Klein Centre for Cosmoparticle Physics. S.V. acknowledges support from the Isaac Newton Trust and the Kavli Foundation through a Newton-Kavli fellowship. We thank the University of Michigan, where part of this work was conducted, for hospitality. L.V. thanks Barry University and the University of Florida for hospitality and support.
\end{acknowledgments}

\bibliographystyle{JHEP}
\bibliography{BHImage}

\end{document}